OpenStudy (anonymous):

f(x)=-x^2+4x-4 four step method f(1)=? f(2)=? f(3)=?

OpenStudy (anonymous):

can someone help me please!!

OpenStudy (anonymous):

r we finding for f(1), f(2) and f(3)?

OpenStudy (ivancsc1996):

Well, you only have to calculate using x=1......\[f(1)=-(1)^{2}+4(1)-4=1\]\[f(2)=-(2)^{2}+4(2)-4=8\]\[f(3)=?\]Try doing f(3) yourself.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that the four step method?

OpenStudy (anonymous):

lol so many steps?

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

as ivancsc1996 showed, it can be done in one

OpenStudy (anonymous):

f'(x)=?

OpenStudy (anonymous):

f'(x)= -2x+4

OpenStudy (ivancsc1996):

Oh that's the derivative which is what HawkCrimson showed.

OpenStudy (anonymous):

f(3)=13?

OpenStudy (ivancsc1996):

No, 17

OpenStudy (anonymous):

i though you subtract 4

OpenStudy (anonymous):

thought

OpenStudy (ivancsc1996):

Oh no it is -1

OpenStudy (anonymous):

f(x)=-x^2+4x-4 f(3)=-(3)^2+4(3)-4 f(3)=-9 +12 -4 f(3) = -1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

find the derivative of (13x^2-26x+26)e^-8x

OpenStudy (anonymous):

u'll have to use the product rule

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so let u = (13x^2-26x+26) and v = e^-8x where U(dv/dx) + V(du/dx)

OpenStudy (anonymous):

13(x^2-2x+2)e^-8x

OpenStudy (anonymous):

thats my answer

OpenStudy (anonymous):

ill do it now but that answer seems too similar to the question, did u use the product rule V(du/dx) + U(dv/dx)?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

im looking at a similar one on my book idk if this is right but its my educated try

OpenStudy (anonymous):

ill tell u the answer in a sec, im almost done

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got something different but i got big numbers lol ill show u my working

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the question is find the derivative with respect to x if y

OpenStudy (anonymous):

so to find dy/dx?

OpenStudy (anonymous):

how i did it was like this y = (13x^2-26x+26)e^-8x and now im solving for dy/dx, isnt that wat u want?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

U = 13x^2 - 26x + 26 V = e^-8x du/dx = 26x - 26 dv/dx = -8e^-8x

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

no, that is half of the working

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

V(du/dx) + U(dv/dx) (e^-8x)(26x - 26) + (13x^2 - 26x + 26)( -8e^-8x) 2e^-8x [ (13x - 13) + -4(13x^2 - 26x + 26) 2e^-8x [ (13x - 13) + (-52x^2 + 104x - 104) ] 2e^-8x [ -52x^2 + 117x - 117 ]

OpenStudy (anonymous):

I dont think its asking for that because i just did a similar problem (15x^3+9x^2+20x-4)^e-5x answer:(3x^3-4x)e^-5x

OpenStudy (anonymous):

I thought it was asking for the opposite way instead of solving

OpenStudy (anonymous):

lol so the question wasnt asking for dy/dx?

OpenStudy (anonymous):

the question is find the derivative with respect to x if y

OpenStudy (anonymous):

whatever that means

OpenStudy (anonymous):

xD so wat method did u use?

OpenStudy (anonymous):

so r u saying my answer is wrong?

OpenStudy (anonymous):

product and chain rule

OpenStudy (anonymous):

idk it just look different than my book

OpenStudy (anonymous):

the answer in my book from the example problem i showed you

OpenStudy (anonymous):

is it (15x^3+9x^2+20x-4)^e-5x or (15x^3+9x^2+20x-4)e^-5x ??